3x - 2y =8 let y = -1
Answer by stanbon(75887)

3x - 2y =8 let y = -1
-----------------
Substitute for "y" and solve for "x":
3x -2(-1) = 8
---
3x + 2 = 8
3x = 6
x = 2
===============
Cheers,
Stan H.